OpenStudy (asiannerd):

Given sin x= -3/5 and x is in quadrant 3, what is the value of tan x/2 ?

MARC (marc_d):

look at the attached file

OpenStudy (mww):

If you are looking for \[\tan \frac{ x }{ 2 }\] then recall this formula \[\sin x = \frac{ 2t }{ 1+t^2 }\] where \[t = \tan(\frac{ x }{ 2 })\] Alternatively find tan(x) using Pythagoras' Theorem:  Then use this formula (half angle t-identity) \[\tan x = \frac{ 2t }{ 1-t^2 }\]

MARC (marc_d):

@asiannerd